# Change-Agent: Towards Interactive Comprehensive Remote Sensing Change Interpretation and Analysis

### v2

@@ IV-D LLM: Brain of Change-Agent
Recently, LLMs have garnered significant attention for their remarkable achievements. By harnessing extensive web knowledge, LLMs have demonstrated substantial promise in attaining human-level intelligence, exhibiting noteworthy capabilities in tasks encompassing instruction comprehension, planning, reasoning, and proficient natural language interactions with humans [58, 59, 60, 61]. Building upon these advancements, there has been a growing research field that employs LLMs as central controllers for constructing autonomous agents [62, 63, 64, 65].
Inspired by these developments, we harness the capabilities of LLMs as the cognitive core responsible for orchestrating the scheduling of our Change-Agent. As depicted in Fig. 8, the LLM meticulously plans task execution based on user instructions. Despite excelling at text-related tasks, LLMs lack inherent visual perception capabilities. To bridge this gap and enable change interpretation and analysis akin to human capability, we equip our agent with a suite of Python tools. These tools include a visual feature extraction backbone, change detection branch, change captioning branch, and relevant Python libraries. Leveraging these tools, the LLM autonomously crafts Python programs, subsequently executed by a Python interpreter to accomplish tasks beyond its intrinsic capabilities. Subsequently, the LLM processes the results and furnishes feedback to the user. In addition to furnishing change masks and descriptions, our agent can implement object counting, image processing, change cause estimation, prediction of future changes, and more. In essence, our Change-Agent adeptly provides users with pertinent information, harnessing the rich knowledge of the LLM in tandem with the provided tools.
@@ V-A Objective Function for Multi-task Training
$\mathcal{L}_{\text{det}}=-\frac{1}{H\times W}\sum_{h=1,w=1}^{H,W}\sum_{j=1}^{C}y_{hw}^{(j)}\log(p_{hw}^{(j)})$ (18)
where $C$ is the number of classes. $H$ and $W$ are the height and width of the change mask. $\bm{p}_{hw}$ = [ $p_{hw}^{(0)},...,p_{cls}^{(C)}$] is the predicted probability vector at the pixel position (h,w)). $\bm{\widetilde{y}}_{hw}$ = [ $\widetilde{y}_{hw}^{(0)},...,\widetilde{y}_{hw}^{(C)}$] is the one-hot vector representation of the ground-truth at the pixel position (h,w)).
@@ V-B Evaluation Metrics
$\text{MIoU}=\frac{1}{\text{N}}\sum_{i=1}^{\text{N}}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}$ (21)
Where N is the total number of classes. For the $i$-th class, $\text{TP}_{i}$, $\text{FP}_{i}$, and $\text{FN}_{i}$ represent the number of true positive, false positive, and false negative, respectively.
@@ V-C Experimental Details
We implemented all models using the PyTorch deep learning framework and trained them on the NVIDIA GTX 4090 GPU. During training, we minimized the loss function defined by Equation 20 and employed the Adam optimizer [71] with an initial learning rate of 0.0001 to optimize the model parameters. We set the maximum number of epochs to 200 and utilized word embeddings with a dimensionality of 512. The backbone for feature extraction is the Siamese weight-shared Segformer-B1 [52]. Initially, we trained the backbone until the sum of BLEU-4 and MIoU indicators did not increase for 50 consecutive epochs. Subsequently, we froze the backbone network and continued training the two branches separately.
@@
Change Detection FC-EF [35] 84.54
FC-Siam-Conc [35] 85.42
FC-Siam-Di [35] 85.20
BIT [46] 84.45
@@ V-D Multi-level Change Interpretation Performance
Currently, there is a dearth of methods that simultaneously address change detection and change captioning. To assess the efficacy of our model in handling both tasks, we conducted comprehensive performance evaluations by comparing against established methods in each respective field on the proposed dataset.
For the change detection evaluation, we have benchmarked several well-known methods, including FC-EF [35], FC-Siam-Conc [35], FC-Siam-Di [35], and BIT [46]. Meanwhile, for the change captioning assessment, we compare our model with several existing state-of-the-art (SOTA) methods, including Capt-Rep-Diff [66], Capt-Att [66], Capt-Dual-Att [66], DUDA [66], MCCFormer-S [72], MCCFormer-D [72], RSICCFormer [13] and Chg2Cap [14].
The comprehensive performance comparisons, as shown in Table II, underscore the superiority of our proposed methodology across both change detection and change captioning tasks. Our method not only achieves simultaneous proficiency in both tasks but also surpasses previous methods. Specifically, our model exhibits superior performance in change detection with a +1.01% improvement on MIoU. Moreover, in terms of change captioning, our model exhibits a noteworthy advancement, showcasing a +1.56% improvement on BLEU-4 and a substantial +3.68% improvement on CIDEr-D.
@@ V-E1 BI3 Layer
As a critical component of our MCI model, the BI3 layer incorporates LPE and GDFA modules. The LPE module enhances feature representation through local perception enhancement,improving the model’s sensitivity to multi-scale changes. The GDFA module effectively integrates the differential information, allowing the model to capture the change area more accurately. In Table III, we conducted ablation experiments on the effectiveness of the LPE and GDFA modules. The baseline utilizes Transformer encoding layers to process bi-temporal features. The experimental results show that integrating the LPE and GDFA modules leads to improved performance in both change detection and change captioning tasks.
Figure 10: The loss curve shows the imbalance in the training process of the two tasks when we add the detection loss and caption loss to form the total loss.
@@ V-F Qualitative Change Interpretation Results
To illustrate the effectiveness of our proposed model in simultaneously generating change masks and change descriptions, Fig. 11 presents comparative results between our proposed model and the baseline model. The experimental results demonstrate the superior capabilities of our model in joint interpretation of change detection and change captioning.
@@ V-G Interaction Between User and Change-Agent
Fig. 13 presents a specific example to illustrate the executable code generated by the LLM within our Change-Agent. When the user requests to perform change detection, display building areas in green, display road areas in blue, and count changed buildings, the LLM efficiently generates executable Python code, demonstrating the agent’s versatility and responsiveness to user needs.
Through these interactive capabilities, our Change-Agent serves as a versatile tool for facilitating nuanced and insightful analysis of changes in remote sensing images, fostering seamless communication and collaboration between users and remote sensing expertise.

### v3

@@ IV-D LLM: Brain of Change-Agent
Recently, LLMs have garnered significant attention for their remarkable achievements. By harnessing extensive web knowledge, LLMs have demonstrated substantial promise in attaining human-level intelligence, exhibiting noteworthy capabilities in tasks encompassing instruction comprehension, planning, reasoning, and proficient natural language interactions with humans [73, 14, 74, 75]. Building upon these advancements, many studies are beginning to utilize LLMs as central controllers for building autonomous agents [76, 77, 78, 79].
Inspired by these developments, we harness the capabilities of LLMs as the cognitive core responsible for orchestrating the scheduling of our Change-Agent. As depicted in Fig. 9, the LLM meticulously plans task execution based on user instructions. Despite excelling at text-related tasks, LLMs lack inherent visual perception capabilities. To bridge this gap and enable change interpretation and analysis akin to human capability, we equip our agent with a suite of Python tools. These tools include a visual feature extraction backbone, change detection branch, change captioning branch, and relevant Python libraries. Leveraging these tools, the LLM autonomously crafts Python programs, subsequently executed by a Python interpreter to automatically accomplish tasks beyond its intrinsic capabilities without the need for human intervention. Subsequently, the LLM processes the results and furnishes feedback to the user. In addition to furnishing change masks and descriptions, our agent can implement object counting, image processing, change cause estimation, prediction of future changes, and more. In short, the Change-Agent adeptly provides users with pertinent information by leveraging the rich knowledge of the LLM and additional tools.
@@ V-A Objective Function for Multi-task Training
$\mathcal{L}_{\text{det}}=-\frac{1}{H\times W}\sum_{h=1,w=1}^{H,W}\sum_{j=1}^{C}\widetilde{y}_{hw}^{(j)}\log(p_{hw}^{(j)})$ (18)
where $C$ is the number of classes. $H$ and $W$ are the height and width of the change mask. $\bm{p}_{hw}$ = [ $p_{hw}^{(0)},...,p_{cls}^{(C)}$] is the predicted probability vector at the pixel position (h,w). $\bm{\widetilde{y}}_{hw}$ = [ $\widetilde{y}_{hw}^{(0)},...,\widetilde{y}_{hw}^{(C)}$] is the one-hot vector representation of the ground-truth at the pixel position (h,w).
@@ V-B Evaluation Metrics
$\text{IoU}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}$ (21)
$\text{MIoU}=\frac{1}{\text{N}}\sum_{i=1}^{\text{N}}\text{IoU}_{i}$ (22)
Where N is the total number of classes (i.e., Background, road, and building). For the $i$-th class, $\text{TP}_{i}$, $\text{FP}_{i}$, and $\text{FN}_{i}$ represent the number of true positive, false positive, and false negative, respectively.
@@ V-C Experimental Details
We implemented all models using the PyTorch deep learning framework and trained them on the NVIDIA GTX 4090 GPU. During training, we minimized the loss function defined by Equation 20 and employed the Adam optimizer with an initial learning rate of 0.0001 to optimize the model parameters. We set the maximum number of epochs to 200 and utilized word embeddings with a dimensionality of 512. The backbone for feature extraction is the Siamese weight-shared Segformer-B1 [56]. In the two branches, we use three BI3 layers. Initially, we trained the backbone until the sum of BLEU-4 and MIoU indicators did not increase for 50 consecutive epochs. Subsequently, we froze the backbone network and continued training the two branches separately.
@@
Change Detection FC-EF [40] 82.70
FC-Siam-Conc [40] 84.25
FC-Siam-Di [40] 84.20
BIT [50] 84.16
ACABFNet [84] 84.43
DARNet [32] 84.99
DMINet [85] 85.37
BiFA [35] 85.68
@@ V-D Multi-level Change Interpretation Performance
Currently, there is a dearth of methods that simultaneously address change detection and change captioning. To assess the efficacy of our model in handling both tasks, we conducted comprehensive performance evaluations by comparing against established methods in each respective field on the proposed LEVIR-MCI dataset.
For the change detection evaluation, we have benchmarked several well-known methods, including FC-EF [40], FC-Siam-Conc [40], FC-Siam-Di [40], BIT [50], ACABFNet [84], DARNet [32], DMINet [85], and BiFA [35]. Meanwhile, for the change captioning assessment, we compare our model with several existing state-of-the-art (SOTA) methods, including Capt-Rep-Diff [64], Capt-Att [64], Capt-Dual-Att [64], DUDA [64], MCCFormer-S [65], MCCFormer-D [65], RSICCFormer [10] and Chg2Cap [11].
Table II shows the comprehensive performance comparisons, underscoring the superiority of our proposed methodology across both change detection and change captioning tasks. Our method not only simultaneously accomplishes two tasks but also surpasses previous single-task methods. Specifically, our model demonstrates leading change detection performance, achieving a +0.75% improvement on MIoU. Additionally, for change captioning, our model exhibits a significant advancement, with a +1.56% improvement on BLEU-4 and a substantial +3.68% improvement on CIDEr-D.
@@ V-E1 BI3 Layer
As a critical component of our MCI model, the BI3 layer incorporates LPE and GDFA modules. The LPE module enhances feature representation through local perception enhancement, improving the model’s sensitivity to multi-scale changes. The GDFA module effectively integrates the differential information, allowing the model to capture the change area more accurately. In Table III, we conducted ablation experiments on the effectiveness of the LPE and GDFA modules. The baseline utilizes Transformer encoding layers to process bi-temporal features. The experimental results show that integrating the LPE and GDFA modules leads to improved performance in two change interpretation tasks, especially in change captioning.
Additionally, considering the various sizes of objects in remote sensing images, it is essential to recognize that small changed objects, due to their fewer pixels, may not contribute significantly to the overall IoU of the test set even if our model accurately detects these small changes. This could obscure the IoU improvement for small objects within the entire test set. Based on this insight, we have added an experiment to further analyse the IoU for small buildings with fewer than 400 pixels. We excluded the IoU analysis of roads as roads with fewer than 400 pixels are very rare. The results, presented in Table VI, show that the IOU for small object change detection is low, indicating the difficulty in detecting small changes. The results also demonstrate that our proposed modules are effective. The qualitative results in Fig. 12 further support this. For instance, in rows 1 and 9 of the figure, our method correctly detects small changed buildings.
@@
TABLE VI: Verification of the effectiveness of the BI3 layer on small objects.
Method GDFA LPE IoUbuilding<400
Baseline 15.52
MCINet-G 15.63
MCINet-L 16.23
Our MCINet 17.24
Figure 11: The loss curve shows the imbalance between tasks during training when we add the detection loss and caption loss to form the total loss.
@@ V-F Qualitative Change Interpretation Results
To illustrate the effectiveness of our proposed model in simultaneously generating change masks and change descriptions, we show the comparison results between our proposed model and the baseline model in Fig. 12. The experimental results demonstrate the superior capabilities of our model in joint interpretation of change detection and change captioning.
@@ V-G Interaction Between User and Change-Agent
Fig. 14 presents a specific example to illustrate the executable code generated by the LLM within our Change-Agent. When the user requests to perform change detection, display building areas in green, display road areas in blue, and count changed buildings, the LLM efficiently generates executable Python code, demonstrating the agent’s versatility and responsiveness to user needs. The execution of the code occurs internally within the Change-Agent. Users simply need to express their requirements without requiring specialized programming skills, which is a significant advantage of our Change-Agent.
Through the interactive capability, our Change-Agent serves as a versatile tool for facilitating nuanced and insightful analysis of changes in remote sensing images, fostering seamless communication and collaboration between users and remote sensing expertise.